Two trains depart simultaneously from a station, traveling in opposite directions. One averages 8 km/h more than the other, and after a 1/2 hour they are 100 km apart. Determine the speed of each train.
Answer by checkley77(12569)

x/2+(x+8)/2=100
(x+x+8)/2=100
2x+8=2*100
2x+8=200
2x=200-8
2x=192
x=192/2
x=96 kmh is the speed of the slower train.
96+8=104 kmh is the speed of the faster train.
Proof:
96/2+104/2=100
48+52=100
100=100